### v1

@@ 1 Introduction
Over the past decade, advancements in model architectures, the availability of larger datasets, and improvements in hardware—among other factors—have significantly enhanced the capabilities of generative machine learning models [1, 2, 3]. At the same time, ongoing progress toward scalable quantum hardware has sparked growing interest in the development of quantum machine learning (QML) algorithms [4, 5], which aim to leverage quantum properties—such as superposition and entanglement—to enhance the efficiency and expressivity of classical machine learning approaches. Although large-scale fault-tolerant quantum hardware is not yet realizable, many QML algorithms are specifically designed to operate within the constraints of the noisy intermediate-scale quantum (NISQ) era [6, 7, 8].
@@
In image generation tasks, several classical deep learning architectures have demonstrated notable effectiveness. Variational Autoencoders (VAEs) are particularly useful for tasks like image denoising [9] and anomaly detection [10] due to their structured latent spaces. Generative Adversarial Networks (GANs) [11] have excelled in realistic image synthesis [12, 13] and image-to-image translation [14, 15], benefiting from their adversarial training mechanism. More recently, diffusion models have emerged as state-of-the-art for tasks such as text-to-image generation [16] and high-resolution image synthesis [17], offering improved output quality and diversity. Quantum analogs of the aforementioned architectures have been proposed [18, 19, 20]; however, their efficiency currently lags behind that of their classical counterparts. It remains an open question whether a fully quantum architecture can effectively perform a generative task while simultaneously demonstrating any form of quantum advantage. To address this, we will focus on quantum versions of Generative Adversarial Networks (QGANs), which have shown the most promising results in image generation among existing quantum approaches.
@@
Classical GANs consist of two competing neural networks—the generator and the discriminator—engaged in a dynamic min-max game. The generator attempts to produce data that mimics the real distribution, while the discriminator aims to distinguish between real and generated data. This adversarial training continues until a Nash equilibrium is reached, ideally when the generator produces outputs indistinguishable from real data. The first proposals toward the development of Quantum Generative Adversarial Networks (QGANs) were introduced in Refs. [19, 21], where the potential advantages of quantum models were demonstrated, while Ref. [21] specifically showed that training is feasible within the quantum framework using a simple circuit. Subsequent studies introduced hybrid QGANs, in which one component—typically the discriminator—is implemented as a classical network, while the other remains quantum. These hybrid models have been applied to learning discrete [22, 23], continuous [24, 23], and random distributions [25]. Image generation tasks have been addressed using hybrid QGANs through various techniques, including a patching scheme [26, 27], learning a discrete distribution [28], employing a quantum Wasserstein GAN architecture [29], and replacing the patching and principal component analysis by a quantum long-short term memory approach [30]. Alongside hybrid approaches, several fully quantum GAN schemes have been proposed. These include entangling QGANs [31], approaches leveraging quantum optimal control techniques [32], and architectures based on the Wasserstein GAN framework [33]. Discrete datasets have also been explored in this context [34], and some of these models have been experimentally implemented on superconducting circuits [35, huang2021quantum]. For image generation tasks QuGAN [36] and IQGAN [37] have shown the most promising results. Similar architectures like EQ-GAN [31] have been used for efficient state preparations, but in principle could be employed for image generation as well.
In this work, we focus on the QuGAN and IQGAN models and evaluate their generalization capabilities. The numerical tests developed for this purpose lead us to explore fundamental limitations of the quantum generator within a QGAN framework.
@@
The paper is organized as follows: In Sec. 2 we evaluate the performance of the state-of-the-art QGAN architectures which is followed by an analytic derivation of a lower bound explaining the failure of generalization of current architectures in Sec. 3. In Sec. 4 we conclude our results. In the appendices A and B we show additional information such as the loss curves and numerical examples, how noise affects a quantum generator.
@@ 2 QGANs for Image Generation
We highlight several key aspects of this approach that, as we will see, are crucial to its performance. First, the model is trained separately on each class of the MNIST dataset. Second, a Principal Component Analysis (PCA) pre-processing step is applied before encoding, which drastically reduces the dimensionality of each training image—from 784 pixels down to just 4. This represents an extreme compression of the data. Additionally, the swap test, which functions as the discriminator in this setup, is not a trainable quantum variational circuit (QVC). By definition, this makes the discriminator non-trainable, distinguishing it from typical adversarial learning frameworks. As a result, the IQGAN circuit aligns more closely with a Quantum Circuit Born Machine (QCBM)[42], acting as an optimized generator without engaging in a min-max adversarial game. Finally, we know that in classical GAN architectures, the generator receives a noise vector as input, enabling the production of varied outputs that resemble the distribution of the training data [11, 43]. In contrast, IQGAN does not incorporate any input noise, which may limit its capacity to generate diverse outputs.
@@
Let us now proceed with the evaluation of the model. In Fig. 1(b)(i), we present the image generated by the model after being trained on class 3 of the MNIST dataset. Note, that the corresponding loss curve is given in Appendix A. Right next to it in Fig. 1(b)(ii), we show the class average for comparison. As can be seen, the two images are nearly identical. This unexpected similarity raised concerns about the actual generative capabilities of the QVC, prompting us to consider whether the PCA step was solely responsible for the observed output. To investigate this further, we replaced the QVC with a random number generator. Surprisingly, the resulting outputs (see Fig. 1(b)(v)) were visually similar—suggesting that the quantum circuit is not contributing to the generation process in its current form.
@@
We have observed that both models exhibit poor generalization performance, despite employing well-motivated architectures and theoretically grounded design choices. In the next section, we aim to shed light on the underlying reasons for this limitation by deriving analytical bounds on the fidelity achieved by QGANs.
@@ 3 Fidelity bounds in QGANs
We want to evaluate the optimal discriminator under the condition $\rho_{\text{data}}=\rho_{G}$, i.e. in the case where the probability density functions of real and generated data coincide. The optimal discriminator for a given generator $G$ is defined by the Helstrom bound [46, 47] that applies in the case of optimal measurement for binary classification:
$D(x)$ $=D(G(z))=\frac{1}{2}+\left\|\pi_{\text{data}}\rho_{\text{data}}-\pi_{\text{G}}\rho_{G}\right\|_{1},$ (2)
where $\|\cdot\|_{1}$ is the trace norm and $\pi_{\text{data}},\ \pi_{\text{G}}=1-\pi_{\text{data}}$ are the prior probabilities corresponding to the class distribution. In the QGAN setting presented, here, both are equally probable and set to $1/2$. Assuming $\rho_{\text{data}}=\rho_{G}$ and inserting Eq. (2) to Eq. (1) we get:
$\min_{G}\max_{D}V(D,G)$ $=\log\left(\frac{1}{2}+\frac{1}{2}\left\|\rho_{\text{data}}-\rho_{G}\right\|_{1}\right)$ (3)
$\quad+\log\left(\frac{1}{2}-\frac{1}{2}\left\|\rho_{\text{data}}-\rho_{G}\right\|_{1}\right)$
$=-2\log 2.$
@@
Using the Helstrom bound, the optimal discriminator can be expressed as a projective measurement onto a specific quantum state $|\delta\rangle$, yielding the following probabilities for real and generated inputs:
$\begin{split}D(x)&=|\langle\delta|x_{i}\rangle|^{2},\\ D(G(z))&=|\langle\delta|\gamma\rangle|^{2},\end{split}$ (5)
where $\langle\delta|=\langle 0|D^{\dagger}$ corresponds to the discriminators measurement, $|\gamma\rangle=G|0\rangle$ is the output state of the generator, and $|x_{i}\rangle=E_{i}|0\rangle$ represents an encoded real data sample. This implies that the discriminator does not depend on individual inputs $x$, but rather on the quantum states $\rho_{\text{data}}$ and $\rho_{G}$.
For any two density matrices $\rho$ and $\sigma$, the trace distance and fidelity are related by the well-known Fuchs–van de Graaf inequality [48]:
@@
$1-F(\rho,\sigma)\leq\frac{1}{2}\left\|\rho-\sigma\right\|_{1}\leq\sqrt{1-F(\rho,\sigma)^{2}}.$ (6)
Applying this to the data and generated quantum states, we obtain the following bound:
$\frac{1}{2}+\frac{1}{2}\left\|\rho_{\text{data}}-\rho_{G}\right\|_{1}\geq 1-\frac{1}{2}F(\rho_{\text{data}},\rho_{G}).$ (7)
Assume that the generated state is pure, $\rho_{G}=|\gamma\rangle\langle\gamma|$. Then the fidelity between the data and generated states is given by:
@@
$F(\rho_{\text{data}},\rho_{G})=\langle\gamma|\rho_{\text{data}}|\gamma\rangle=\sum_{i}p_{\text{data},i}|\langle\gamma|x_{i}\rangle|^{2},$ (8)
where we have also used $\rho_{\text{data}}=\sum_{i=1}^{r}p_{data,i}|x_{i}\rangle\langle x_{i}|$, where r corresponds to the number of data points. The best achievable fidelity is obtained when $|\gamma\rangle=|x_{\mathrm{max}}\rangle$, where $|x_{\mathrm{max}}\rangle=\mathrm{argmax}_{i}(\langle x_{i}|\rho_{\text{data}}|x_{i}\rangle)$.
Therefore according to Eq. (7) the best-case fidelity a pure-state generator can achieve is bounded:
$\frac{1}{2}+\frac{1}{2}\left\|\rho_{\text{data}}-\rho_{G}\right\|_{1}$ $\geq 1-\frac{p_{\text{data},\text{max}}}{2}.$ (9)
@@
In Fig. 3, we present the absolute value of the state $|x_{\mathrm{max}}\rangle$ for each class of the MNIST dataset. These visualizations closely resemble $\left|x_{\text{max}}\right\rangle$ and are similar to the optimal outputs generated by both QuGAN (see Fig. 2) and IQGAN (see Fig. 1), supporting our analytical finding that these represent the optimal outputs achievable by the models. The more complex (i.e., higher-rank) a dataset is, the poorer the training performance of a QGAN with a pure-state generator, due to fundamental limits on fidelity and distinguishability. This limitation extends to other quantum generative models that output a single pure quantum state. In such cases, if no sampling is performed, the task shifts from generalizing an unknown distribution to simply fitting or reproducing a known one. While this may be appropriate for applications like efficiently representing quantum states [31] or implementing quantum look-up tables [49], it does not constitute a generalizing approach.
To this end, it is meaningful to explore the generalization capabilities of quantum generators that do not necessarily output pure states. A discussion of this direction, along with some preliminary results, is provided in Appendix B.

### v2

@@ 1 Introduction
Over the past decade, advancements in model architectures, the availability of larger datasets, and improvements in hardware—among other factors—have significantly enhanced the capabilities of generative machine learning models [1, 2, 3]. At the same time, ongoing progress toward scalable quantum hardware has sparked growing interest in the development of quantum machine learning (QML) algorithms [4, 5], which aim to leverage quantum properties—such as superposition and entanglement—to enhance the efficiency and expressivity of classical machine learning approaches. Although large-scale fault-tolerant quantum hardware is not yet realizable, many QML algorithms are specifically designed to operate within the constraints of the noisy intermediate-scale quantum (NISQ) era [6, 7, 8]. Within this broader context, QML has been applied to quantum-state analysis, including entanglement characterization, as well as to variational models for quantum and classical data, highlighting its growing scope in near-term quantum systems. [9, 10, 11, 12, 13].
@@
In image generation tasks, several classical deep learning architectures have demonstrated notable effectiveness. Variational Autoencoders (VAEs) are particularly useful for tasks like image denoising [14] and anomaly detection [15] due to their structured latent spaces. Generative Adversarial Networks (GANs) [16] have excelled in realistic image synthesis [17, 18] and image-to-image translation [19, 20], benefiting from their adversarial training mechanism. More recently, diffusion models have emerged as state-of-the-art for tasks such as text-to-image generation [21] and high-resolution image synthesis [22], offering improved output quality and diversity.
Quantum analogues of the aforementioned architectures have been proposed [23, 24, 25]; however, their efficiency currently lags behind that of their classical counterparts. In particular, it is still unclear whether fully quantum generative models can genuinely learn complex data distributions rather than reproducing dominant features of the training data. In this work, we therefore focus on Quantum Generative Adversarial Networks (QGANs), which have reported the most promising results for quantum image generation to date.
@@
Classical GANs consist of two competing neural networks—the generator and the discriminator—engaged in a dynamic min-max game. The generator attempts to produce data that mimics the real distribution, while the discriminator aims to distinguish between real and generated data. This adversarial training continues until a Nash equilibrium is reached, ideally when the generator produces outputs indistinguishable from real data. The first proposals toward the development of Quantum Generative Adversarial Networks (QGANs) were introduced in Refs. [24, 26], where the potential advantages of quantum models were demonstrated, while Ref. [26] specifically showed that training is feasible within the quantum framework using a simple circuit. Subsequent studies introduced hybrid QGANs, in which one component—typically the discriminator—is implemented as a classical network, while the other remains quantum. These hybrid models have been applied to learning discrete [27, 28], continuous [29, 28], and random distributions [30]. Image generation tasks have been addressed using hybrid QGANs through various techniques, including a patching scheme [31, 32], learning a discrete distribution [33], employing a quantum Wasserstein GAN architecture [34], and replacing the patching and principal component analysis by a quantum long-short term memory approach [35]. Alongside hybrid approaches, several fully quantum GAN schemes have been proposed. These include entangling QGANs [36], approaches leveraging quantum optimal control techniques [37], and architectures based on the Wasserstein GAN framework [38]. Discrete datasets have also been explored in this context [39], and some of these models have been experimentally implemented on superconducting circuits [40, 41]. For image generation tasks QuGAN [42] and IQGAN [43] have shown the most promising results. Similar architectures like EQ-GAN [36] have been used for efficient state preparations, but in principle could be employed for image generation as well.
In this work, we focus on the QuGAN and IQGAN models and evaluate their generalization capabilities. Specifically, we investigate whether their reported performance reflects genuine distribution learning or instead convergence to dominant features of the training data. Our numerical findings motivate a closer analysis of the generator’s role within the QGAN framework, ultimately revealing fundamental limitations that arise when the generator outputs a single pure quantum state.
@@
The paper is organized as follows: In section ”QGANs for Image Generation” we evaluate the performance of the state-of-the-art QGAN architectures which is followed by an analytic derivation of a lower bound explaining the failure of generalization of current architectures in section ”Fidelity Bounds in QGANs”. In section ”Conclusions” we conclude our results. In the Appendices we show additional information such as the loss curves and numerical examples, how noise affects a quantum generator.
@@ 2 QGANs for Image Generation
At this point, it is worth clarifying the distinction between adversarial learning and a Quantum Circuit Born Machine (QCBM). A QCBM optimizes a single parametrized quantum circuit using a direct, often fidelity-based, objective rather than an adversarial discriminator. As a result, the generator is trained in isolation instead of through a min–max game, leading to fundamentally different training dynamics. This distinction becomes relevant when interpreting the behavior of IQGAN below.
We highlight several key aspects of this approach that, as we will see, are crucial to its performance. First, the model is trained separately on each class of the MNIST dataset. Second, a Principal Component Analysis (PCA) pre-processing step is applied before encoding, which drastically reduces the dimensionality of each training image—from 784 pixels down to just 4. This represents an extreme compression of the data. Additionally, the swap test, which functions as the discriminator in this setup, is not a trainable quantum variational circuit (QVC). By definition, this makes the discriminator non-trainable, distinguishing it from typical adversarial learning frameworks. The swap test therefore induces a fixed, fidelity-based cost function rather than a learnable adversarial objective. As a result, the IQGAN circuit aligns more closely with a Quantum Circuit Born Machine (QCBM)[48], acting as an optimized generator without engaging in a min-max adversarial game. Finally, we know that in classical GAN architectures, the generator receives a noise vector as input, enabling the production of varied outputs that resemble the distribution of the training data [16, 49]. In contrast, IQGAN does not incorporate any input noise, which may limit its capacity to generate diverse outputs.
@@
Let us now proceed with the evaluation of the model. In Fig. 1(b)(i), we present the image generated by the model after being trained on class 3 of the MNIST dataset. Note, that the corresponding loss curve is given in the “Supplementary Material”. Right next to it in Fig. 1(b)(ii), we show the class average for comparison. As can be seen, the two images are nearly identical. This unexpected similarity raised concerns about the actual generative capabilities of the QVC, prompting us to consider whether the PCA step was solely responsible for the observed output. To investigate this further, we replaced the QVC with a random number generator. Surprisingly, the resulting outputs (see Fig. 1(b)(v)) were visually similar—suggesting that the quantum circuit is not contributing to the generation process in its current form.
@@
To complement the qualitative analysis presented above, we now turn to a quantitative evaluation of the generated samples using standard distribution-level metrics [52, 53]. Specifically, we assess the similarity between real and quantum-generated images by computing the Fréchet Inception Distance [54] (FID) on MNIST (digit 3) and CIFAR-10 (class 0), which we select as representative benchmark classes. After training, the quantum generator is sampled by performing projective measurements on the generator qubits in the computational basis, yielding binary bitstrings that serve as latent representations. For example, with four generator qubits, a single measurement may produce a bitstring such as $0101$, corresponding to a four-dimensional latent vector. Repeating this procedure yields an empirical distribution over the corresponding latent representations.
These bitstrings are mapped back to the continuous data space by applying the inverse MinMax normalization followed by the inverse principal component analysis (PCA), resulting in reconstructed grayscale images. Both real and generated images are resized to $299\times 299$, replicated across three channels, and processed using a pretrained Inception-v3 network with its classification head removed to extract 2048-dimensional feature representations. The mean $\mu$ and covariance $\Sigma$ of these features are computed for each dataset, and the FID is defined as
$\mathrm{FID}=\|\mu_{\mathrm{real}}-\mu_{\mathrm{gen}}\|_{2}^{2}+\mathrm{Tr}\!\left(\Sigma_{\mathrm{real}}+\Sigma_{\mathrm{gen}}-2\bigl(\Sigma_{\mathrm{real}}\Sigma_{\mathrm{gen}}\bigr)^{1/2}\right),$
where lower values indicate greater similarity between the real and generated data distributions.
In addition to samples obtained from the trained quantum generators, we evaluate a classical reference baseline based on uniformly random sampling in the latent space. Specifically, for a fixed number of generator qubits, we sample binary bitstrings uniformly at random, corresponding to projective measurements of an unstructured generator that assigns equal probability to all computational basis states. These randomly sampled bitstrings are treated identically to the quantum-generated samples. They are mapped to continuous latent vectors via inverse MinMax normalization, transformed back to the data space using the inverse PCA, and post-processed into images that are subsequently used for FID evaluation. This random-sampling baseline represents the absence of any learned structure in the latent distribution and provides a reference for evaluating whether adversarial training improves the quality of the generated samples.
Figure 3: FID values as a function of the number of principal components. For each model, samples drawn from the trained generator circuit (blue markers with solid lines) are compared against a classical baseline obtained by uniformly random sampling of latent bitstrings (orange markers with dashed lines).
As shown in Fig. 3, across different numbers of PCA components, both IQGAN and QUGAN achieve FID values that are comparable to those obtained from the uniformly random baseline. This behavior indicates that, within the considered regimes, adversarial training fails to induce a meaningful latent distribution that improves upon unstructured random sampling. The near overlap between the trained-generator curves and the random baseline thus provides clear evidence of the limited expressive power of the investigated quantum circuits at the tested system sizes and circuit depths. Notably, in the IQGAN setting applied to the more complex CIFAR-10 dataset, we observe regimes in which uniform random sampling yields lower FID values than sampling from the trained quantum generator, further underscoring that adversarial training does not reliably improve sample quality.
We have observed that both models exhibit poor generalization performance, despite employing well-motivated architectures and theoretically grounded design choices. A crucial point we wish to emphasize is that, in both IQGAN and QuGAN, there is no sampling over latent variables, no classical mixing over circuit parameters, and no stochastic measurement-induced distribution. Consequently, the generator represents the entire model distribution by a single deterministic quantum state. To ensure reproducibility, in Appendix B we provide information on the training procedures and hyperparameters used in our numerical simulations. Motivated by the aforementioned observation, in the next section, we shed light on the underlying reasons for this limitation by deriving analytical bounds on the fidelity achieved by QGANs.
@@ 3 Fidelity bounds in QGANs
For fixed generator state $\rho_{G}$ and data state $\rho_{\text{data}}$, the optimal minimum-error binary measurement for discriminating between the two states, given prior probabilities $\pi_{\text{data}}$ and $\pi_{G}$, attains the Helstrom success probability [55, 56]
$P_{\mathrm{succ}}^{\star}=\frac{1}{2}\left(1+\left\|\pi_{\text{data}}\rho_{\text{data}}-\pi_{G}\rho_{G}\right\|_{1}\right),$ (2)
where $\|\cdot\|_{1}$ denotes the trace norm and $\pi_{G}=1-\pi_{\text{data}}$. In the balanced setting $\pi_{\text{data}}=\pi_{G}=\tfrac{1}{2}$ this becomes $P_{\mathrm{succ}}^{\star}=\tfrac{1}{2}\left(1+\tfrac{1}{2}\|\rho_{\text{data}}-\rho_{G}\|_{1}\right)$.
Identifying the discriminator’s optimal average output with this success probability, i.e. $D^{\star}=D(x)=D(G(z)):=P_{\mathrm{succ}}^{\star}$, we obtain for the value function
$V(D^{\star},G)$ $=\log\!\left(\frac{1}{2}+\frac{1}{4}\left\|\rho_{\text{data}}-\rho_{G}\right\|_{1}\right)+\log\!\left(\frac{1}{2}-\frac{1}{4}\left\|\rho_{\text{data}}-\rho_{G}\right\|_{1}\right).$ (3)
In particular, at $\rho_{\text{data}}=\rho_{G}$ we have $\|\rho_{\text{data}}-\rho_{G}\|_{1}=0$ and therefore
$V(D^{\star},G)=\log\!\left(\tfrac{1}{2}\right)+\log\!\left(\tfrac{1}{2}\right)=-2\log 2.$ (4)
@@
In the quantum setting, the discriminator corresponds to a fixed measurement optimized to distinguish whether an input state originates from the data ensemble $\rho_{\text{data}}$ or from the generator ensemble $\rho_{G}$. For given $\rho_{\text{data}}$ and $\rho_{G}$, the optimal strategy is the Helstrom measurement [55, 56], defined by the positive and negative eigenspaces of the operator $\pi_{\text{data}}\rho_{\text{data}}-\pi_{G}\rho_{G}$. In the balanced case, this measurement depends only on the difference $\rho_{\text{data}}-\rho_{G}$. For concreteness, and without loss of generality in the pure-state case, the Helstrom measurement may be represented as a projective measurement onto the eigenbasis of this operator. Let $|\delta\rangle$ denote a normalized eigenstate associated with the positive eigenspace. The discriminator output probabilities for real and generated inputs are then
$\begin{split}D(x_{i})&=|\langle\delta|x_{i}\rangle|^{2},\\ D(G(z))&=|\langle\delta|\gamma\rangle|^{2},\end{split}$ (6)
where $|x_{i}\rangle=E_{i}|0\rangle$ denotes an encoded real data sample drawn from the ensemble $\rho_{\text{data}}=\sum_{i}p_{i}|x_{i}\rangle\langle x_{i}|$, and $|\gamma\rangle=G|0\rangle$ is the generator output corresponding to latent input $z$. Although the measurement outcome probabilities depend on the particular input state, the discriminator itself is fully specified by a single fixed measurement determined solely by $\rho_{\text{data}}$ and $\rho_{G}$, and does not adapt to individual samples. In this sense, the quantum discriminator is directly analogous to the classical optimal discriminator, which is completely determined by the underlying data and generator distributions.
The distinguishability of $\rho_{\text{data}}$ and $\rho_{G}$ appearing in the Helstrom bound can be related to the fidelity via the Fuchs–van de Graaf inequalities [57],
$1-F(\rho_{\text{data}},\rho_{G})\leq\frac{1}{2}\left\|\rho_{\text{data}}-\rho_{G}\right\|_{1}\leq\sqrt{1-F(\rho_{\text{data}},\rho_{G})^{2}},$ (7)
providing a direct connection between optimal discrimination performance and fidelity-based metrics commonly used in quantum generative modeling. Applying this inequality yields
$\frac{1}{2}+\frac{1}{4}\left\|\rho_{\text{data}}-\rho_{G}\right\|_{1}\geq 1-\frac{1}{2}F(\rho_{\text{data}},\rho_{G}),$ (8)
@@
where by $F(\rho,\sigma)=\operatorname{Tr}\left(\sqrt{\sqrt{\rho}\,\sigma\,\sqrt{\rho}}\right)$ we denote the fidelity.
If the generated state is pure, $\rho_{G}=|\gamma\rangle\langle\gamma|$, then
$F(\rho_{\text{data}},\rho_{G})=\sqrt{\langle\gamma|\rho_{\text{data}}|\gamma\rangle}.$ (9)
@@
For an ensemble decomposition $\rho_{\text{data}}=\sum_{i=1}^{r}p_{\text{data},i}\,|x_{i}\rangle\langle x_{i}|$, this overlap evaluates to
$\langle\gamma|\rho_{\text{data}}|\gamma\rangle=\sum_{i=1}^{r}p_{\text{data},i}\,|\langle\gamma|x_{i}\rangle|^{2}.$ (10)
The maximum overlap is achieved by choosing $|\gamma\rangle$ to be an eigenvector of $\rho_{\text{data}}$ corresponding to its largest eigenvalue $\lambda_{\max}(\rho_{\text{data}})$, yielding
$F_{\max}=\sqrt{\lambda_{\max}(\rho_{\text{data}})}.$ (11)
Consequently, Eq. (8) implies the following best-case bound for any pure-state generator:
$\frac{1}{2}+\frac{1}{4}\left\|\rho_{\text{data}}-\rho_{G}\right\|_{1}\geq 1-\frac{1}{2}\sqrt{\lambda_{\max}(\rho_{\text{data}})}.$ (12)
The bound in Eq. (12) depends on the data state only through its largest eigenvalue $\lambda_{\max}(\rho_{\text{data}})$, and is therefore governed by the spectral concentration of the data ensemble. If $\rho_{\text{data}}$ has rank $r$, then $\lambda_{\max}(\rho_{\text{data}})\geq 1/r$, yielding the worst–case scaling $F_{\max}\sim 1/\sqrt{r}$ for flat spectra. For low–rank or spectrally peaked data, $\lambda_{\max}(\rho_{\text{data}})$ is close to unity. Equality in Eq. (12) is attained in the extremal cases where $\rho_{\text{data}}=\rho_{G}$, corresponding to vanishing trace distance, or when $\rho_{\text{data}}$ and $\rho_{G}$ have orthogonal support, in which case $\|\rho_{\text{data}}-\rho_{G}\|_{1}=2$ and the fidelity becomes zero. For mixed data states, a pure-state generator cannot achieve vanishing trace distance, instead, the optimal solution corresponds to recovering the principal eigenstate of $\rho_{data}$, which maximizes the fidelity while remaining at finite trace distance from the full data ensemble.
In Fig. 4, we present the absolute value of the principal eigenstate $|x_{\max}\rangle$ of the data density operator $\rho_{\text{data}}$ for each class of the MNIST dataset, where $|x_{\max}\rangle$ denotes the eigenvector associated with the largest eigenvalue $\lambda_{\max}(\rho_{\text{data}})$. Importantly, this state does not in general coincide with any individual data sample, but instead represents a coherent superposition capturing the dominant mode of the data ensemble. The close visual resemblance between $|x_{\max}\rangle$ and the outputs generated by QuGAN (Fig. 2) and IQGAN (Fig. 1) indicates that both models converge toward this principal eigenstate. This behavior is consistent with our analytical results, which show that for generators restricted to producing a single pure quantum state, the maximal achievable fidelity with $\rho_{\text{data}}$ is attained by the leading eigenvector of the data density operator. The observed convergence therefore reflects the optimal pure-state approximation of the data ensemble, rather than successful learning of the full data distribution.
@@
The more complex (i.e., higher-rank) a dataset is, the poorer the training performance of a QGAN with a pure-state generator, due to fundamental limits on fidelity and distinguishability. This limitation extends to other quantum generative models that output a single pure quantum state. In such cases, if no sampling is performed, the task shifts from generalizing an unknown distribution to simply fitting or reproducing a known one. While this may be appropriate for applications like efficiently representing quantum states [36] or implementing quantum look-up tables [58], it does not constitute a generalizing approach. To this end, it is meaningful to explore the generalization capabilities of quantum generators that do not necessarily output pure states. A discussion of this direction, along with some preliminary results, is provided in Appendix C.